OpenStudy (anonymous):

Ifthegrowthrateofthenumber of bacteria at any time t is proportional to the number present at t and doubles in 1 week, how many bacteria can be expected after 2 weeks? After 4 weeks?

OpenStudy (ilovecake):

et x be the number of bacteria. The growth rate is the change in x over time dxdt growth rate proportional to the number present means dxdt=k⋅x where k is the constant of proportionality

OpenStudy (ilovecake):

oops

OpenStudy (anonymous):

i was able to get to that part \[\frac{ dx }{ dt }=k*x\]

OpenStudy (aum):

\[\large \frac{dx}{x} = k*dt \\ \large \ln(x) = kt \\ \large x = x_0e^{kt} \]

OpenStudy (aum):

where \(x_0\) is the number of bacteria at time t = 0. When t = 1, \(x = 2x_0\): \(2x_0 = x_0e^k;~~~~2 = e^k; ~~~~k = \ln(2)\)

OpenStudy (anonymous):

so then id place 2x0=e^ln(2)?

OpenStudy (aum):

\[ x = x_0e^{kt} \\ k = \ln(2) \\ x = x_0e^{\ln(2) * t} = x_0(e^{\ln(2)})^t = x_02^t \\ \text{When t = 2, }x = x_02^2 = 4x_0 ~~\text{ (4 times the initial population in 2 weeks.) } \\ \text{When t = 4, }x = x_02^4 = 16x_0 ~~\text{ (16 times the initial population in 4 weeks.) } \\ \]

OpenStudy (anonymous):

ohhh okay!

OpenStudy (aum):

\(e^{\ln(2)} = 2 \)

OpenStudy (anonymous):

sweet thanks aum!